OpenStudy (anonymous):

a circle has a diameter of 24 cm. what is its radius?

OpenStudy (anonymous):

its 12

OpenStudy (anonymous):

see its diameter if we take r=24 then radius would be r/2=24/2=12 cm

OpenStudy (anonymous):

Thank you <3